In a random sample of 25 residents of the state of New Hampshire, the mean waste recycled per person per day was 1.0 pounds with a standard deviation of 0.58 pounds. Determine the 98% confidence interval for the mean waste recycled per person per day for the population of New Hampshire. Assume the population is approximately normal. Step 1 of 2 : Find the critical value that should be used in constructing the confidence interval. Round your answer to three decimal places.
Answer #1

Solution :

Given that,

\bar x = 1.0

s = 0.58

n = 25

Degrees of freedom = df = n - 1 = 25 - 1 = 24

At 98% confidence level the t is ,

\alpha = 1 - 98% = 1 - 0.98 = 0.02

\alpha / 2 = 0.02 / 2 = 0.01

t\alpha /2  df = t0.01,24= 2.492

critical value is 2.492

Margin of error = E = t\alpha/2,df * (s /\sqrtn)

= 2.492 * (0.58 / \sqrt 25) = 0.289

The 98% confidence interval estimate of the population mean is,

\bar x - E < \mu < \bar x + E

1.0 - 0.289 < \mu < 1.0+ 0.289

0.711 < \mu < 1.289

(0.711 , 1.289)